Bernoulli's DE

(2y^3-x^3)dx+3x^2ydy=0
1
Expert's answer
2021-07-15T06:17:37-0400

Solution;

Rewrtie the equation as follows:

"\frac{dy}{dx}" +P(x)y=Q(x)yn

"\frac{2y^3-x^3}{3x^2y}+\frac{dy}{dx}=0"

"\frac{dy}{dx}-\frac{x}{3y}=-\frac{2}{3x^2}y^2"

The equation is not a Bernoulli's so we solve as an homogenous equation.

"\frac{dy}{dx}" =-"\frac{2y^3-x^3}{3x^2y}"

Take

y=Vx

"\frac{dy}{dx}" =v+x"\frac{dv}{dx}"

v+x"\frac{dv}{dx}" =-"\frac{2v^3x^3-x3}{3vx^3}"

Divide the R.H.S with x3

v+x"\frac{dv}{dx}" ="\frac{1-2v^3}{3v}"

Separate the variables;

"\frac{3v}{1-2v^3-3v^2}"dv="\frac1x" dx

Integrating both sides;

3("-\frac{ln(2v-1)}{9}" +"\frac{ln(v+1)}{9}"+"\frac{3}{3v+3}" =ln(x)+C

But v="\frac yx"

-"\frac{ln(\frac yx-1)}{3}" +"\frac{ln(\frac yx+1)}{3}"+"\frac{3}{3\frac yx+3}" =ln(x)+C